OpenStudy (anonymous):

fid the derivative: y=e^sin^2(lnx)

OpenStudy (ipwnbunnies):

\[y = e^{\sin^{2} (lnx)}\]?

OpenStudy (anonymous):

yes

OpenStudy (ipwnbunnies):

Let's use the chain rule" \[y = e^{u} \frac{du}{dx}\]

OpenStudy (ipwnbunnies):

dy/dx = ...*

OpenStudy (ipwnbunnies):

So yeah, let's find the derivative of the power, first.

OpenStudy (ipwnbunnies):

\[\sin^{2}(lnx)\] Some more chain rule here.

OpenStudy (anonymous):

u= lnx ?

OpenStudy (ipwnbunnies):

Uh, yeah, for this part.

OpenStudy (anonymous):

u= 1/x

OpenStudy (anonymous):

\[e^{\sin ^{2}\ln x} 2\sin( \ln x) \cos (\ln x)(1/x)\]

OpenStudy (anonymous):

Is this the right solution?

OpenStudy (ipwnbunnies):

I believe so. You see, we kept the e^u part, but we multiply it by the derivative of u to get the derivative of the entire thing.

OpenStudy (anonymous):

soso707 then what is the right solution?????

OpenStudy (ipwnbunnies):

Many chain rules.

OpenStudy (anonymous):

but I don't understand? can you understand me >>?

OpenStudy (anonymous):

I want steps for this Question

OpenStudy (ipwnbunnies):

Ok. Let's go back to the general form: \[\frac{dy}{dx} = e^{u} \frac{du}{dx}\]

OpenStudy (ipwnbunnies):

As you see, the derivative of an exponential function with e involves us keep the original function, but multiplying is by the derivative of its power.

OpenStudy (ipwnbunnies):

So, we're doing this: \[\frac{dy}{dx} = e^{\sin^{2}(lnx)}* \frac{d(\sin^2 (lnx))}{dx}\]

OpenStudy (anonymous):

ok, I think I'm understand

OpenStudy (ipwnbunnies):

Great. Can you find the derivative of the second part, now?

OpenStudy (anonymous):

which second part?

OpenStudy (mathmale):

\[\frac{dy}{dx} = e^{\sin^{2}(lnx)}* \frac{d(\sin^2 (lnx))}{dx}\]

OpenStudy (mathmale):

The "second" or "latter" part is \[ \frac{d(\sin^2 (lnx))}{dx}\]

OpenStudy (mathmale):

Please try that. Remember the power rule? Use that first.

OpenStudy (anonymous):

(sin^2 . 1/x) + (2cos . lnx) ?

OpenStudy (mathmale):

would you mind doing this easier problem first?\[\frac{ d }{ dx }(2\cos ^{2}x)\]

OpenStudy (mathmale):

I'm asking you to do this for practice in using the power and chain rules.

OpenStudy (anonymous):

-4sinx

OpenStudy (mathmale):

Here's how I'd approach that example problem: 1. Take the coefficient (2) out: \[2\frac{ d }{ dx }(\cos ^{2}x)\]

OpenStudy (anonymous):

ok ..

OpenStudy (mathmale):

2. Think of \[\cos ^{2}x. as. (\cos x)^2\]

OpenStudy (mathmale):

Find the derivative of \[(\cos x)^{2}\]

OpenStudy (mathmale):

Remember to use the power rule and chain rules here.

OpenStudy (anonymous):

2(cosx) -sinx if you write like this, that true ?

OpenStudy (mathmale):

You're on the right track. There's a set of parentheses missing in your 2(cosx) -sinx. Can you figure out where the parentheses are missing and insert them, please? The derivative of (cos x)^2 is :

OpenStudy (anonymous):

2(cosx)(-sinx) like that ?

OpenStudy (mathmale):

so much better. That's perfect. Without the 2nd set of parentheses, it'd appear that you were subtracting sin x instead of multiplying by (-sin x). OK. Now let me change the problem a bit by replacing x with a function: find the derivative of (cos ln x)^2. Remember to use the power rule first, and then the Chain Rule, and to use those lovely parentheses!

OpenStudy (anonymous):

2(cos ln x) (cos . 1/x + (-sin . ln x))

OpenStudy (mathmale):

\[\frac{ d }{ dx }(\cos \ln x)^{2}=2(\cos \ln x)^{1}\frac{ d }{ dx }\ln x\]

OpenStudy (mathmale):

First you take the derivative of (the power function (cos ln x)^2, which you have done correctly; the correct result is 2(cos ln x). But after that you must apply the chain rule and find the derivative of ln x:

OpenStudy (anonymous):

2(cos lnx)^1 (1/x) ?

OpenStudy (mathmale):

Really nice. You're right on target. What have you learned here?

OpenStudy (ipwnbunnies):

Err, what about the -sin(lnx)

OpenStudy (mathmale):

Err, I overlooked that, and you've overlooked the necessary parentheses again. Mind going back and fixing up your result to include the derivative of cos ln x?

OpenStudy (anonymous):

I learned 2 rules

OpenStudy (mathmale):

Your initial response was "2(cos ln x) (cos . 1/x + (-sin . ln x))" Can you fix that?

OpenStudy (mathmale):

First you'll want to write that = (cos ln x), followed by the derivative of cos ln x, followed by the derivative of ln x. Try that , please.

OpenStudy (anonymous):

2(cos ln x) (1/x) ?

OpenStudy (mathmale):

See my previous comment.

OpenStudy (mathmale):

As a review, here's what we have so far in this EXAMPLE problem:\[\frac{ d }{ dx }(\cos \ln x)^2=2(\cos \ln x)^1( ? )( ? ))\]

OpenStudy (anonymous):

2(cos ln x)(cos lnx)(1/x)

OpenStudy (mathmale):

Your 2(cos ln x)(cos lnx)(1/x) is on the right track, but you haven't yet found the derivative of cos ln x, have you? thus, throw out your (cos lnx) term (in the middle) and replace that with the derivative of (cos ln x).

OpenStudy (anonymous):

can you the true sulotion for this 2(cos ln x) (cos . 1/x + (-sin . ln x)) ?

OpenStudy (anonymous):

write *

OpenStudy (mathmale):

I'll take what you have: 2(cos ln x)(cos lnx)(1/x) and write out a corrected version: 2(cos ln x)(-sin lnx)(1/x). That's it. Note that this problem we've been working on is an EXAMPLE; we have not found the derivative of the original problem.

OpenStudy (mathmale):

Original problem: Find the derivative of y=e^sin^2(lnx), or \[y = e^{\sin^{2} (lnx)}\]

OpenStudy (mathmale):

First, what is the derivative of e^x?

OpenStudy (anonymous):

xe

OpenStudy (mathmale):

Hi, Winner66! Of course. Go right ahead!

OpenStudy (mathmale):

@soso707: Would you please look up "derivative of exponential functions" . Winner66 is presenting an alternative way of finding the derivative, a good and valid method. However, I'd like to finish the original approach before considering an alternative approach. What is the derivative of e^x?

OpenStudy (mathmale):

@loser66: Your method will, of course, succeed, but I think it would be helpful to @soso707 if he and I were to finish applying the original approach. Once that's clear, we could consider your suggested "logarithmic differentiation." Thanks.

OpenStudy (mathmale):

First, what is the derivative of e^x? Answer: e^x. Second, what is the derivative of e^(2x)? @soso707 ?

OpenStudy (anonymous):

yes thanks :)

OpenStudy (mathmale):

Second, what is the derivative of e^(2x)? @soso707 ?

OpenStudy (mathmale):

We need the Chain Rule here. the derivative of e^x is simply e^x. The derivative of e^u, where u is a separate function of x, is e^u*(du/dx). Have you seen this before?

OpenStudy (mathmale):

\[\frac{ d }{dx }e ^{u}=e ^{u}*\frac{ du }{ dx }\]

OpenStudy (mathmale):

You will see this again and again in Calculus problems.

OpenStudy (anonymous):

ok if I have e^2x = e^2 ?

OpenStudy (mathmale):

therefore, the derivative of e^(2x) is\[\frac{ e }{ dx}e ^{2x}=e ^{2x }*\frac{ d }{ dx}(2x)=?\]

OpenStudy (mathmale):

Your tentative answer was partially correct, but you didn't apply the chain rule (to find the derivative of 2x).

OpenStudy (anonymous):

can you give me some example ?

OpenStudy (mathmale):

Sure. \[\frac{ d }{ dx }e ^{2x}=e ^{2x}*\frac{ d }{ dx }(2x)=e ^{2x}2 = 2e ^{2x}\]

OpenStudy (mathmale):

Now, follow that example closely and find the derivative of e^(3x).

OpenStudy (mathmale):

the derivative of e^x is simply e^x, or (e^x)*1; the derivative of e^(2x) is e^(2x)*2; the derivative of e^(3x) is ????

OpenStudy (anonymous):

e^(3x)*3 ?

OpenStudy (mathmale):

Very nice. Thanks. Now, try finding the derivative of e^(sin x), following the same rules. Hint: in the end you must differentiate sin x.

OpenStudy (anonymous):

e^(sinx)(cosx) ?

OpenStudy (mathmale):

Perfect! You've gotten it!

OpenStudy (mathmale):

what is the derivative of \[e ^{\sin ^{2}x}?\]follow the same rules as before. But now you must also apply the power rule, as well as the exponential function, trig function and chain rules.

OpenStudy (mathmale):

BRB (be right back)

OpenStudy (anonymous):

e^sin2x(2cos^1x) ?

OpenStudy (mathmale):

Did you find the derivative of (sin x)^2? Remember the power rule with chain rule.

OpenStudy (anonymous):

or e^sin2x 2sinx ?

OpenStudy (mathmale):

Yes, that's a good start; you must now find the derivative of sin x and multiply your previous result by that derivative.

OpenStudy (anonymous):

e^sin2x (2sinx)

OpenStudy (mathmale):

\[\frac{ d }{ dx }e ^{\sin ^{2}x}=e ^{\sin ^{2}x}(2sinx)\frac{ d }{ dx }\sin x=?\]

OpenStudy (mathmale):

So your result is correct, except that y ou have not yet found the derivative of sin x (which is required by the chain rule).

OpenStudy (anonymous):

e^sin2x (2sinx)(cosx)

OpenStudy (mathmale):

that's great. You're on your way.

OpenStudy (anonymous):

thanks, so y=e^sin2(lnx) e^sin2(lnx) 2sin(lnx) cos(lnx) (1/x) ?

OpenStudy (mathmale):

Except that you should label your answer with "(dy/dx)", that's great. Congrats!

OpenStudy (anonymous):

dy/dx= e^sin2(lnx) 2sin(lnx) cos(lnx) (1/x) ?

OpenStudy (mathmale):

Great. Whew!

OpenStudy (anonymous):

yes yes :) :) thanks a lot :)

OpenStudy (mathmale):

I'd bet you're anxious to move on to other problems or other things. But @loser66 had a good point: if you were to take the natural log of both sides of the original, given equation, you'd get\[\ln y = (\sin \ln x)^2\]and finding the derivative (dy/dx) of that would be faster and possibly involve fewer steps than those we've gone through.

OpenStudy (mathmale):

It's up to you: move on to some other problem, or explore this alternative approach.